# Green functions, Hitchin’s formula and curvature equations on tori

### v1

@@ Abstract.
Let $G(z)=G(z;\tau)$ be the Green function on the torus $E_{\tau}=\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ with the singularity at 0. Lin and Wang (Ann. Math. 2010) proved that $G(z)$ has either $3$ or $5$ critical points (depending on the choice of $\tau$). Later, Bergweiler and Eremenko (Proc. Amer. Math. Soc. 2016) gave a new proof of this remarkable result by using anti-holomorphic dynamics.
In this paper, firstly, we prove that once $G(z)$ has $5$ critical points, then these $5$ critical points are all non-degenerate. Secondly, we prove that for any $p\in E_{\tau}$ satisfying $p\neq-p$ in $E_{\tau}$, the number of critical points of $G_{p}(z):=\frac{1}{2}(G(z+p)+G(z-p))$ belongs to $\{4,6,8,10\}$ (depending on the choice of $(\tau,p)$) and each number really occurs. We apply Hitchin’s formula to prove the generic non-degeneracy of critical points. This allows us to study the distribution of the numbers of critical points of $G_{p}$ as $p$ varies. Applications to the curvature equation $\Delta u+e^{u}=4\pi(\delta_{p}+\delta_{-p})$ on $E_{\tau}$ are also given. How the geometry of the torus affects the solution structure is studied.
@@ 1. Introduction
In this paper, we study the Green function of the torus. One of our motivations for this research is to understand how the geometry of the torus affects the solution structure of elliptic PDEs. Consider the curvature equation
@@
    There is a sequence of even solutions $u_{p}(z)$ of (1.1) such that $u_{p}(z)$ blows up as $0\neq p\to 0$ if and only if $G(z;\tau)$ has exactly $3$ critical points and one of them is degenerate.
@@
    as $0\neq p\to 0$ if and only if $G(z;\tau)$ has exactly $5$ critical points.
@@
    There is small $\varepsilon>0$ such that (1.1) has no solutions for any $0<|p|<\varepsilon$ if and only if $G(z;\tau)$ has exactly $3$ critical points that are all non-degenerate.
@@ 1.1. The Green function $G$
The Green function $G(z)=G(z;\tau)$ of the torus $E_{\tau}$ is defined by
$\Delta G(z;\tau)=\delta_{0}-\frac{1}{\left|E_{\tau}\right|}\text{ \ on }E_{\tau},\quad\int_{E_{\tau}}G(z;\tau)=0,$
where $\Delta=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}$ is the Laplace operator, $\delta_{0}$ is the Dirac measure at 0 and $|E_{\tau}|$ is the area of the torus $E_{\tau}$. It is an even function on $E_{\tau}$ with the only singularity at 0. For $k\in\{1,2,3\}$, since
@@
Therefore, nontrivial critical points, if exist, must appear in pairs. Lin and Wang [18] studied the number of critical points of $G(z;\tau)$, and proved the following remarkable result.
@@
A consequence of Theorems B and 1.2 is that if $G(z)$ has a degenerate critical point, then $G(z)$ has only trivial critical points, and this degenerate critical point is trivial. It is well known that the non-degeneray of critical points of Green functions has many important applications, such as in constructing bubbling solutions via the reduction method and proving the local uniqueness of bubbling solutions for elliptic PDEs; see e.g. [2, 9, 20, 21] and references therein. We believe that Theorem 1.2 will have important applications to other problems.
@@ 1.2. The Green function $G_{p}$
The study of (1.1) and the proof of Theorem 1.1 lead us to study the Green fuction $G_{p}(z)$, which is defined by
@@
In Section 2, we will follow the approach of using anti-holomorphic dynamics from Bergweiler-Eremenko [5] to prove Theorem 1.3. After Theorem 1.3, a natural question is whether there exists $(\tau,p)$ such that $G_{p}(z)=G_{p}(z;\tau)$ has exactly $N$ critical points for every $N\in\{4,6,8,10\}$. Recently, it was proved in [12] that the lower bound $4$ occurs; we will recall this result in Theorem 3.5. Our next result confirms that every $N\in\{6,8,10\}$ also occurs.
@@
    There exist $\tau$ such that $G_{\frac{\omega_{k}}{4}}(z;\tau)$ has exactly $6$ critical points with $\det D^{2}G_{\frac{\omega_{k}}{4}}$ $(\frac{\omega_{l}}{2};\tau)<0$ for all $l\in\{0,1,2,3\}$, but the unique pair of nontrivial critical points are degenerate minimal points of $G_{\frac{\omega_{k}}{4}}(z;\tau)$.
@@
    There exist $\tau$ such that $G_{\frac{\omega_{k}}{4}}(z;\tau)$ has exactly $6$ critical points, among which there are exactly $2$ trivial critical points being degenerate, but the other $4$ critical points (including the pair of nontrivial critical points) are all non-degenerate saddle points of $G_{\frac{\omega_{k}}{4}}(z;\tau)$. Furthermore,
@@
(1.8) $\wp(p;\tau)=\wp(r+s\tau;\tau)+\frac{\wp^{\prime}(r+s\tau;\tau)}{2(\zeta(r+s\tau;\tau)-r\eta_{1}(\tau)-s\eta_{2}(\tau))}.$
It is well known that $\wp(\cdot;\tau):E_{\tau}\to\mathbb{C}\cup\{\infty\}$ is a double cover with branch points at $\frac{\omega_{k}}{2}$’s, i.e. for any $c\in\mathbb{C}\cup\{\infty\}$, there is a unique pair $\pm z_{c}\in E_{\tau}$ such that $\wp(\pm z_{c})=c$. Consequently, for any $(r,s)\in\mathbb{R}^{2}\backslash\frac{1}{2}\mathbb{Z}^{2}$, there exists a unique pair $\pm p=\pm p_{r,s}(\tau)\in E_{\tau}$ such that (1.8) holds. Note that $p_{r,s}(\tau)\in E_{\tau}[2]$ might hold for some $(r,s)\in\mathbb{R}^{2}\backslash\frac{1}{2}\mathbb{Z}^{2}$. Therefore, (1.8) implies that there is at most one pair $\pm p\in E_{\tau}\setminus E_{\tau}[2]$ such that $q=r+s\tau$ is a nontrivial critical point of $G_{p}$. This argument also implies the following result.
@@
Then Theorem 1.7 says that $((\mathcal{S}(\wp(p))\cap\mathbb{R}^{2})/\mathbb{Z}^{2})\leq 6$. In general, one might ask whether there is an upper bound for $((\mathcal{S}(\wp(p))\cap((t_{1},t_{2})+\mathbb{R}^{2}))/\mathbb{Z}^{2})$ for given $t_{1},t_{2}\in\mathbb{C}$. We will prove in Remark 2.9 that the upper bound is $10$ for any $t_{1},t_{2}\in\mathbb{C}$.
Conversely, we apply Hitchin’s formula (1.8) to prove the non-degeneracy of all critical points for generic $p$.
@@
that is, $\mathcal{B}_{k}$ is either an open disk or an open half plane. We will prove in Section 3 that $|\alpha_{k}|=\beta_{k}$ is equivalent to that $\frac{\omega_{k}}{2}$ is a degenerate critical point of $G(z)$. It was proved in [13, Theorem 1.2] (we will recall it in Theorem 3.6) that for any $\tau$, there is at most one trivial critical point of $G(z;\tau)$ being degenerate. This implies that there is at most one half plane among $\mathcal{B}_{k}$’s. For example, if $\tau=e^{\pi i/3}$ or $\tau=i$, then $\mathcal{B}_{k}$’s are all disks, the figures of which are presented in Figures 1 and 2. Our next result gives the criterion for the degeneracy of trivial critical points.
@@
Figure 1. The four disks and $\{e_{1},e_{2},e_{3}\}$ for $\tau=e^{\pi i/3}$: the smallest circle for $\partial\mathcal{B}_{0}$, left for $\partial\mathcal{B}_{1}$, upper right for $\partial\mathcal{B}_{2}$ and lower right for $\partial\mathcal{B}_{3}$.
@@
Clearly the number of critical points changes when $\wp(p)\in\Xi$ crosses $\partial\Omega_{6}\cap\partial\Omega_{10}$. Theorem 1.12 shows that Theorem 1.10-(1) is sharp. Theorems 1.10 and 1.12 will be proved in Section 5, where we will also apply Theorem 1.10 to two special cases $\tau=e^{\pi i/3}$ and $\tau=i$; See Theorems 5.7 and 5.8 for details.
Finally, concerning the degeneracy of nontrivial critical points, Theorem 1.4-(2) indicates that nontrivial critical points could be degenerate for some $(\tau,p)$. Here we propose the following conjecture.
@@ 1.3. The curvature equation
For convenience we use the complex notation $z=x+iy$, so the Laplace operator $\Delta=4\partial^{2}_{\bar{z}z}$. By changing variable $z\mapsto z+\frac{p_{1}+p_{2}}{2}$, we can always assume $p_{2}=-p_{1}$, namely we only need to study (1.1):
@@
[12] Let $p\in E_{\tau}\setminus E_{\tau}[2]$. Then the followings hold.
@@ 2. Possible numbers of critical points
The function $g$ has at most $4$ attracting fixed points, modulo $\Lambda_{\tau}$.
@@
The set of critical points of $\widetilde{h}$ belongs to at most $8$ grand orbits under $\widetilde{h}$.
@@
If any of $\{\pm c_{1},\pm c_{2}\}$ is in $\widetilde{F}$, then it is a zero of $\widetilde{h}^{\prime}$. By (2.8), we see that the other zeros of $\widetilde{h}^{\prime}$ are $\widetilde{g}^{-1}(\pm c_{j})$. Note that any point in $\widetilde{g}^{-1}(\pm c_{j})$ belongs to the same grand orbit as $\widetilde{g}(\pm c_{j})$. Therefore, the critical points of $\widetilde{h}$ are contained in at most $8$ grand orbits represented by $\pm c_{j},\widetilde{g}(\pm c_{j})$, $j=1,2$. ∎
@@
Lemma 2.6 implies that if $\widetilde{h}^{n}(\pm c_{j})\to z_{0}$ for some fixed point $z_{0}$ of $\widetilde{g}$, then $\widetilde{h}^{n}(\widetilde{g}(\pm c_{j}))\to z_{0}$. Clearly attracting fixed points of $\widetilde{g}$ are always attracting fixed points of $\widetilde{h}=\widetilde{g}^{2}$, so Fatou’s theorem and Lemma 2.6 together imply that each attracting fixed point of $\widetilde{g}$ attracts at least two critical orbits of $\widetilde{h}$. Then by applying Fatou’s theorem to $\widetilde{h}:S\to S$ (where $S$ is any connected component of $\widetilde{F}$ that contains an attracting fixed point of $\widetilde{g}$) and Lemma 2.5, we deduce that $\widetilde{g}:\widetilde{F}\to\widetilde{F}$ has at most $4$ attracting fixed points
Since the $\pi$-image of an attracting fixed point of $g$ is an attracting fixed point of $\widetilde{g}$, we obtain that $g$ has at most $4$ attracting fixed points, modulo $\Lambda_{\tau}$.
@@
and $\overline{\partial}g_{w}(z)\equiv\overline{\partial}g(z)$. Therefore, by applying the same arguments to $g_{w}$, we conclude that $g_{w}$ has at most $4$ attracting fixed points, modulo $\Lambda_{\tau}$.
@@
By (2.4), we know that the map $\phi:E_{\tau}\to\overline{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$ , $z\mapsto\phi(z):=z-g(z)$ is well-defined. Notice that the set of fixed points of $g$ is precisely $\phi^{-1}(0)$, so our goal is to prove $N:=\#\phi^{-1}(0)\leq 10$.
@@
Since $J_{\phi}=1-|\overline{\partial}g|^{2}>0$ is equivalent to $|\overline{\partial}g|<1$, we see that the zeros of $\phi$ in $D^{+}$ are precisely the attracting fixed points of $g$, so Lemma 2.2 implies $N^{+}\leq 4$. Thus $N=2N^{+}+2\leq 10$.
Step 2. For the general case that 0 might be a singular value of $\phi$, we can modify the proof as follows. Let $w\in\mathbb{C}$ be any regular value of $\phi$, i.e. $\phi^{-1}(w)\subset D^{+}\cup D^{-}$. Note that $\phi(z)=w$ is equivalent to $z=g(z)+w=g_{w}(z)$, so $\phi^{-1}(w)$ is precisely the set of fixed points of $g_{w}$. Thanks to Remark 2.7, we can repeat the same argument as Step 1 to obtain that $\phi^{-1}(w)\leq 10$. Therefore, $\phi^{-1}(w)\leq 10$ for any regular value $w$ of $\phi$.
@@
Since Sard theorem says that the set of singular values of $\phi$ is of Lebesgue measure 0, we conclude from Lemma 2.8 that $\phi^{-1}(w)\leq 10$ for any $w\in\mathbb{C}$ and in particular, $\phi^{-1}(0)\leq 10$.
This proves that $G_{p}(z)$ has at most $10$ critical points. Since $\frac{\omega_{k}}{2}$ are all critical points of $G_{p}$ for $k\in\{0,1,2,3\}$, and nontrivial critical points must appear in pairs, we conclude that the number of critical points of $G_{p}(z)$ belongs to $\{4,6,8,10\}$. The proof is complete. ∎

### v2

@@ Abstract.
Let $G(z)=G(z;\tau)$ be the Green function on the flat torus $E_{\tau}=\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ with the singularity at 0. Lin and Wang (Ann. Math. 2010) proved that $G(z)$ has either $3$ or $5$ critical points (depending on the choice of $\tau$). Later, Bergweiler and Eremenko (Proc. Amer. Math. Soc. 2016) gave a new proof of this remarkable result by using anti-holomorphic dynamics.
In this paper, firstly, we prove that once $G(z)$ has $5$ critical points, then these $5$ critical points are all non-degenerate. Secondly, we study the sum of two Green functions which can be reduced to $G_{p}(z):=\frac{1}{2}(G(z+p)+G(z-p))$. We prove that for any $p$ satisfying $p\neq-p$ in $E_{\tau}$, the number of critical points of $G_{p}(z)$ belongs to $\{4,6,8,10\}$ (depending on the choice of $(\tau,p)$) and each number really occurs. We apply Hitchin’s formula (J. Differ. Geom. 1995) in a surprising way to prove the generic non-degeneracy of critical points. This allows us to study the distribution of the numbers of critical points of $G_{p}(z)$ as $p$ varies. Applications to the curvature equation $\Delta u+e^{u}=4\pi(\delta_{p}+\delta_{-p})$ on $E_{\tau}$ are also given, and how the geometry of the torus affects the solution structure is studied.
@@ 1. Introduction
In this paper, we study the Green function of the flat torus. One of our motivations for this research is to understand how the geometry of the torus affects the solution structure of elliptic PDEs. Consider the curvature equation
@@
    There is a sequence of even solutions $u_{p}(z)$ of (1.1) such that $u_{p}(z)$ blows up as $0\neq p\to 0$ if and only if $G(z)$ has exactly $3$ critical points and one of them is degenerate.
@@
    as $0\neq p\to 0$ if and only if $G(z)$ has exactly $5$ critical points.
@@
    There is small $\varepsilon>0$ such that (1.1) has no solutions for any $0<|p|<\varepsilon$ if and only if $G(z)$ has exactly $3$ critical points that are all non-degenerate.
@@ 1.1. The Green function $G$
The Green function $G(z,w)=G(z,w;\tau)$ of the flat torus $E_{\tau}$ is the unique function that satisfies
$\Delta_{z}G(z,w)=\delta_{w}-\frac{1}{\left|E_{\tau}\right|}\text{ \ on }E_{\tau},\quad\int_{E_{\tau}}G(z,w)dxdy=0,$
where we use the complex variable $z=x+iy$, $\Delta_{z}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}=4\partial^{2}_{\bar{z}z}$ is the Laplace operator, $\delta_{w}$ is the Dirac measure at $w$ and $|E_{\tau}|$ is the area of the torus $E_{\tau}$. By the translation invariance of $\Delta_{z}$, we have G(z,w)=G(zw,0)G(z,w)=G(z-w,0) and it is enough to consider the Green function $G(z;\tau)=G(z):=G(z,0)$. Clearly $G(z)$ is an even function on $E_{\tau}$ with the only singularity at 0, so a-a is also a critical point of $G(z)$ if $a\in E_{\tau}\setminus\{0\}$ is. For $k\in\{1,2,3\}$, since
@@
Therefore, nontrivial critical points, if exist, must appear in pairs. Lin and Wang [18] studied the number of critical points of $G(z)$, and proved the following remarkable result.
@@
A consequence of Theorems B and 1.2 is that if $G(z)$ has a degenerate critical point, then $G(z)$ has only trivial critical points. It is well known that the non-degeneray of critical points of Green functions has many important applications, such as in constructing bubbling solutions via the reduction method and proving the local uniqueness of bubbling solutions for elliptic PDEs; see e.g. [2, 9, 20, 21] and references therein. We believe that Theorem 1.2 will have important applications to other problems.
@@ 1.2. The Green function $G_{p}$
The study of (1.1) and the proof of Theorem 1.1 lead us to study the sum of two Green fuctions $G(z-p_{1})+G(z-p_{2})$. By changing variable $z\mapsto z+\frac{p_{1}+p_{2}}{2}$, we can always assume $p_{2}=-p_{1}$, so it is enough to study the Green function $G_{p}(z)=G_{-p}(z)$, which is defined by
@@
Moreover, if $\wp(2p)=2\eta_{1}-\frac{2\pi}{\operatorname{Im}\tau}$, then $G_{p}(z)$ has at most one pair of nontrivial critical points, or equivalently, the number of critical points of $G_{p}(z)$ belongs to $\{4,6\}$.
Here $\wp(z)$ and $\eta_{1}=\eta_{1}(\tau)$ are classical special functions that will be recalled soon. In Section 2, we will follow the approach of using anti-holomorphic dynamics from Bergweiler-Eremenko [5] to prove Theorem 1.3. After Theorem 1.3, a natural question is whether there exists $(\tau,p)$ such that $G_{p}(z)=G_{p}(z;\tau)$ has exactly $N$ critical points for every $N\in\{4,6,8,10\}$. This question can not be settled from the proof of Theorem 1.3. Recently, it was proved in [12] that the lower bound $4$ occurs; we will recall this result in Theorem 3.5. By developing new ideas, our next result confirms that every $N\in\{6,8,10\}$ also occurs.
@@
    There exists $\tau$ such that $G_{\frac{\omega_{k}}{4}}(z;\tau)$ has exactly $6$ critical points with $\det D^{2}G_{\frac{\omega_{k}}{4}}$ $(\frac{\omega_{l}}{2};\tau)<0$ for all $l\in\{0,1,2,3\}$, but the unique pair of nontrivial critical points are degenerate minimal points of $G_{\frac{\omega_{k}}{4}}(z;\tau)$.
@@
    There exists $\tau$ such that $G_{\frac{\omega_{k}}{4}}(z;\tau)$ has exactly $6$ critical points, among which there are exactly $2$ trivial critical points being degenerate, but the other $4$ critical points (including the pair of nontrivial critical points) are all non-degenerate saddle points of $G_{\frac{\omega_{k}}{4}}(z;\tau)$. Furthermore,
@@
(1.8) $\wp(p)=\wp(r+s\tau)+\frac{\wp^{\prime}(r+s\tau)}{2(\zeta(r+s\tau)-r\eta_{1}-s\eta_{2})}.$
It is well known that $\wp(\cdot):E_{\tau}\to\mathbb{C}\cup\{\infty\}$ is a double cover with branch points at $\frac{\omega_{k}}{2}$’s, i.e. for any $c\in\mathbb{C}\cup\{\infty\}$, there is a unique pair $\pm z_{c}\in E_{\tau}$ such that $\wp(\pm z_{c})=c$. Consequently, for any $(r,s)\in\mathbb{R}^{2}\backslash\frac{1}{2}\mathbb{Z}^{2}$, there exists a unique pair $\pm p=\pm p_{r,s}(\tau)\in E_{\tau}$ such that (1.8) holds. Note that $p_{r,s}(\tau)\in E_{\tau}[2]$ might hold for some $(r,s)\in\mathbb{R}^{2}\backslash\frac{1}{2}\mathbb{Z}^{2}$. Therefore, (1.8) implies that there is at most one pair $\pm p\in E_{\tau}\setminus E_{\tau}[2]$ such that $q=r+s\tau$ is a nontrivial critical point of $G_{p}$. This argument also implies the following result.
@@
Note that $(-r,-s)\in\mathcal{S}(\wp(p))$ if $(r,s)\in\mathcal{S}(\wp(p))$. Then Theorem 1.7 says that $((\mathcal{S}(\wp(p))\cap\mathbb{R}^{2})/\mathbb{Z}^{2})\leq 6$. In general, one might ask whether there is an upper bound for $((\mathcal{S}(\wp(p))\cap((t_{1},t_{2})+\mathbb{R}^{2}))/\mathbb{Z}^{2})$ for given $t_{1},t_{2}\in\mathbb{C}$. We will prove in Remark 2.9 that the upper bound is $10$ for any $t_{1},t_{2}\in\mathbb{C}$.
Conversely, we apply Hitchin’s formula (1.8) in a quite different way to prove the non-degeneracy of all critical points for generic $p$.
@@
that is, $\mathcal{B}_{k}$ is either an open disk or an open half plane. We will prove in Section 3 that $|\alpha_{k}|=\beta_{k}$ is equivalent to that $\frac{\omega_{k}}{2}$ is a degenerate critical point of $G(z)$. It was proved in [13, Theorem 1.2] (we will recall it in Theorem 3.7) that for any $\tau$, there is at most one trivial critical point of $G(z;\tau)$ being degenerate. This implies that there is at most one half plane among $\mathcal{B}_{k}$’s. For example, if $\tau=e^{\pi i/3}=\frac{1}{2}+\frac{\sqrt{3}}{2}i$ or $\tau=i$, then $\mathcal{B}_{k}$’s are all disks, the figures of which are presented in Figures 1 and 2. Our next result gives the criterion for the degeneracy of trivial critical points.
@@
Figure 1. The four disks and $\{e_{1},e_{2},e_{3}\}$ for $\tau=e^{\pi i/3}=\frac{1}{2}+\frac{\sqrt{3}}{2}i$: the smallest circle for $\partial\mathcal{B}_{0}$, left for $\partial\mathcal{B}_{1}$, upper right for $\partial\mathcal{B}_{2}$ and lower right for $\partial\mathcal{B}_{3}$.
@@
Clearly the number of critical points changes when $\wp(p)\in\Xi$ crosses $\partial\Omega_{6}\cap\partial\Omega_{10}$. Theorem 1.12 shows that Theorem 1.10-(1) is sharp. Theorems 1.10 and 1.12 will be proved in Section 5, where we will also apply Theorem 1.10 to two special cases $\tau=\frac{1}{2}+\frac{\sqrt{3}}{2}i$ and $\tau=i$; See Theorems 5.7 and 5.8 for details. In subsequet work, we will continue to study $G_{p}(z)$ for $\tau=ib$ or $\tau=\frac{1}{2}+ib$ with $b>0$, i.e. when $E_{\tau}$ is a general rectangular torus or rhombus torus.
Finally, concerning the degeneracy of nontrivial critical points, Theorem 1.4-(2) and Theorem 1.12 indicate that nontrivial critical points could be degenerate for some $(\tau,p)$. Here we propose the following conjecture.
@@ 1.3. The curvature equation
By changing variable $z\mapsto z+\frac{p_{1}+p_{2}}{2}$, we can always assume $p_{2}=-p_{1}$, namely we only need to study (1.1):
@@
[12] Let $p\in E_{\tau}\setminus E_{\tau}[2]$. Then the following statements hold.
@@ 2. Possible numbers of critical points
The function $g$ has at most $4$ attracting fixed points (resp. at most $2$ attracting fixed points if $\wp(2p)=-2a=2\eta_{1}-\frac{2\pi}{\operatorname{Im}\tau}$), modulo $\Lambda_{\tau}$.
@@
The set of critical points of $\widetilde{h}$ belongs to at most $8$ grand orbits (resp. at most $4$ grand orbits if $\wp(2p)=-2a$) under $\widetilde{h}$.
@@
If any of $\{\pm c_{1},\pm c_{2}\}$ is in $\widetilde{F}$, then it is a zero of $\widetilde{h}^{\prime}$. By (2.8), we see that the other zeros of $\widetilde{h}^{\prime}$ are $\widetilde{g}^{-1}(\pm c_{j})$. Note that any point in $\widetilde{g}^{-1}(\pm c_{j})$ belongs to the same grand orbit as $\widetilde{g}(\pm c_{j})$. Therefore, the critical points of $\widetilde{h}$ are contained in at most $8$ grand orbits represented by $\pm c_{j},\widetilde{g}(\pm c_{j})$, $j=1,2$.
Now we further assume $\wp(2p)=-2a$. Then by the additional formula of elliptic functions
$\wp(u-v)=-\wp(u)-\wp(v)+\frac{(\wp^{\prime}(u)+\wp^{\prime}(v))^{2}}{4(\wp(u)-\wp(v))^{2}},$
we see that $\wp(z+p)+\wp(z-p)=2a=-\wp(2p)$ implies $\wp^{\prime}(z+p)+\wp^{\prime}(z-p)=0,$ so $\wp(z+p)+\wp(z-p)=2a$ has only $2$ solutions $\pm c_{1}=\pm c_{2}$ of both multiplicities $2$ on $E_{\tau}$. Consequently, the critical points of $\widetilde{h}$ are contained in at most $4$ grand orbits represented by $\pm c_{1},\widetilde{g}(\pm c_{1})$. ∎
@@
Lemma 2.6 implies that if $\widetilde{h}^{n}(\pm c_{j})\to z_{0}$ for some fixed point $z_{0}$ of $\widetilde{g}$, then $\widetilde{h}^{n}(\widetilde{g}(\pm c_{j}))\to z_{0}$. Clearly attracting fixed points of $\widetilde{g}$ are always attracting fixed points of $\widetilde{h}=\widetilde{g}^{2}$, so Fatou’s theorem and Lemma 2.6 together imply that each attracting fixed point of $\widetilde{g}$ attracts at least two critical orbits of $\widetilde{h}$. Then by applying Fatou’s theorem to $\widetilde{h}:S\to S$ (where $S$ is any connected component of $\widetilde{F}$ that contains an attracting fixed point of $\widetilde{g}$) and Lemma 2.5, we deduce that $\widetilde{g}:\widetilde{F}\to\widetilde{F}$ has at most $4$ attracting fixed points (resp. at most $2$ attracting fixed points if $\wp(2p)=-2a$).
Since the $\pi$-image of an attracting fixed point of $g$ is an attracting fixed point of $\widetilde{g}$, we obtain that $g$ has at most $4$ attracting fixed points (resp. at most $2$ attracting fixed points if $\wp(2p)=-2a$), modulo $\Lambda_{\tau}$.
@@
and $\overline{\partial}g_{w}(z)\equiv\overline{\partial}g(z)$. Therefore, by applying the same arguments to $g_{w}$, we conclude that $g_{w}$ has at most $4$ attracting fixed points (resp. at most $2$ attracting fixed points if $\wp(2p)=-2a$), modulo $\Lambda_{\tau}$.
@@
By (2.4), we know that the map $\phi:E_{\tau}\to\overline{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, $z\mapsto\phi(z):=z-g(z)$ is well-defined. Notice that the set of fixed points of $g$ is precisely $\phi^{-1}(0)$, so our goal is to prove $N:=\#\phi^{-1}(0)\leq 10$ (resp. $N\leq 6$ if $\wp(2p)=-2a$).
@@
Since $J_{\phi}=1-|\overline{\partial}g|^{2}>0$ is equivalent to $|\overline{\partial}g|<1$, we see that the zeros of $\phi$ in $D^{+}$ are precisely the attracting fixed points of $g$, so Lemma 2.2 implies $N^{+}\leq 4$ (resp. $N^{+}\leq 2$ if $\wp(2p)=-2a$). Thus $N=2N^{+}+2\leq 10$ (resp. $N\leq 6$ if $\wp(2p)=-2a$).
Step 2. For the general case that 0 might be a singular value of $\phi$, we can modify the proof as follows. Let $w\in\mathbb{C}$ be any regular value of $\phi$, i.e. $\phi^{-1}(w)\subset D^{+}\cup D^{-}$. Note that $\phi(z)=w$ is equivalent to $z=g(z)+w=g_{w}(z)$, so $\phi^{-1}(w)$ is precisely the set of fixed points of $g_{w}$. Thanks to Remark 2.7, we can repeat the same argument as Step 1 to obtain that $\phi^{-1}(w)\leq 10$ (resp. $\phi^{-1}(w)\leq 6$ if $\wp(2p)=-2a$). Therefore, $\phi^{-1}(w)\leq 10$ (resp. $\phi^{-1}(w)\leq 6$ if $\wp(2p)=-2a$) for any regular value $w$ of $\phi$.
@@
Since Sard theorem says that the set of singular values of $\phi$ is of Lebesgue measure 0, we conclude from Lemma 2.8 that $\phi^{-1}(w)\leq 10$ (resp. $\phi^{-1}(w)\leq 6$ if $\wp(2p)=-2a$) for any $w\in\mathbb{C}$ and in particular, $\phi^{-1}(0)\leq 10$ (resp. $\phi^{-1}(0)\leq 6$ if $\wp(2p)=-2a$).
This proves that $G_{p}(z)$ has at most $10$ critical points (at most $6$ if $\wp(2p)=-2a$). Since $\frac{\omega_{k}}{2}$ are all critical points of $G_{p}$ for $k\in\{0,1,2,3\}$, and nontrivial critical points must appear in pairs, we conclude that the number of critical points of $G_{p}(z)$ belongs to $\{4,6,8,10\}$ (resp. $\{4,6\}$ if $\wp(2p)=-2a$). The proof is complete. ∎